OpenStudy (anonymous):

Why does this (-2)^2 equal 4 and not -4?

OpenStudy (anonymous):

squares of numbers are always positive !!

OpenStudy (johnweldon1993):

exactly...but remember if we had -(2)^2....what would THAT be?

OpenStudy (anonymous):

negative x negative equals positive. so -2 x -2 =4 and not -4

OpenStudy (anonymous):

\[(-n)^2=(-1)^2*n^2\]

OpenStudy (anonymous):

and -1*-1=1

OpenStudy (anonymous):

that will be 4 because the negative sign is not being squared

OpenStudy (johnweldon1993):

it's important to remember the parenthesis.... if we have (-2)^2 ...that = 4 if we have -(2)^2 ...that = -4...the - comes in AFTER the fact....the 2 gets squared...then the answer is turned negative

OpenStudy (anonymous):

here's a tip: If the exponent is even number, the answer is positive. And if the exponent is odd number, the answer will be negative

OpenStudy (amistre64):

(-2)^2 = -2 * -2 - (2 * -2) = -(-4) = 4

OpenStudy (anonymous):

Thank you all very much! I was confused.

OpenStudy (johnweldon1993):

no problem...hope you understand it better now